How many different letter arrangements are possible using all the letters in the word "evaluate"?
I thought I had to do something like
8 / 2!2! = 8 / 4 = 2
But on my test I got it wrong and I knew it couldn't be the right answer. This subject really confuses me!

Answer by stanbon(75887)

How many different letter arrangements are possible using all the letters in the word "evaluate"?
I thought I had to do something like
8 / 2!2! = 8 / 4 = 2
But on my test I got it wrong
-------------
Correct Answer:: 8!/(2!*2!) = 40,320/4 = 10,080
===============
Cheers,
Stan H.
----------------